\begin{thebibliography}{12}


\bibitem{racahmatrices}
G. Racah, %Theory of Complex Spectra. II.
Phys.Rev. {\bf 62} (1942) 438-462\\
E.P. Wigner, Manuscript, 1940,  in: {\sl Quantum Theory of Angular Momentum},
pp. 87–133, Acad.Press, 1965;
{\sl Group Theory and Its Application to the Quantum Mechanics of Atomic Spectra},
Acad.Press,   1959\\
L.D. Landau and E.M. Lifshitz, {\sl Quantum Mechanics: Non-Relativistic Theory},
Pergamon Press, 1977 \\
J. Scott Carter, D.E. Flath, M. Saito, {\sl The Classical and Quantum 6j-symbols},
Princeton Univ.Press, 1995 \\
S. Nawata, P. Ramadevi and Zodinmawia, Lett.Math.Phys. {\bf 103} (2013) 1389-1398,
arXiv:1302.5143\\
A. Mironov, A. Morozov, A. Sleptsov,  	JHEP 07 (2015) 069, arXiv:1412.8432


\bibitem{modtrans}
D. Galakhov, A. Mironov, and A. Morozov,
%S-Duality and Modular Transformation as a non-perturbative deformation of the ordinary pq-duality,
JHEP 06 (2014) 050, arXiv/1311.7069\\
N.Nemkov,  Eur.Phys.J. C77 (2017) 368,  arXiv:1610.02000 \\
M. Billo, M. Frau, F. Fucito, A. Lerda, J. F. Morales, R. Poghossian, D. Ricci Pacifici,
%Modular anomaly equations in N = 2  theories and their large-N limit.
JHEP 10 (2014) 131, arXiv:1406.7255 \\
G.Aminov, A.Mironov and A.Morozov, JHEP (2017),  arXiv:1709.04897

\bibitem{CFT}  A. Belavin, A. Polyakov, and A. Zamolodchikov,
%"Infinite Conformal Symmetry in Two-Dimensional Quantum Field Theory,"
Nucl.Phys. B241 (1984) 333–380\\
A.Zamolodchikov, Al.Zamolodchikov,
{\it Conformal field theory and critical phenomena in 2d systems}, 2009\\
%, 168 p. \\
L.Alday, D.Gaiotto and Y.Tachikawa, Lett.Math.Phys. 91 (2010) 167-197, arXiv:0906.3219

\bibitem{knotinvs}
 J.W.Alexander, Trans.Amer.Math.Soc. 30 (2) (1928) 275-306\\
V.F.R.Jones, Invent.Math. 72 (1983) 1 Bull.AMS 12 (1985) 103 Ann.Math. 126 (1987) 335\\
L.Kauffman, Topology 26 (1987) 395\\
P.Freyd, D.Yetter, J.Hoste, W.B.R.Lickorish, K.Millet, A.Ocneanu, Bull. AMS. 12 (1985) 239\\
J.H.Przytycki and K.P.Traczyk, Kobe J Math. 4 (1987) 115-139\\
A.Morozov, Theor.Math.Phys. 187 (2016) 447-454, arXiv:1509.04928

\bibitem{CS}
S. Chern and J. Simons, Ann.Math. 99 (1974) 48-69\\
E. Witten, Comm.Math.Phys. 121 (1989) 351-399

\bibitem{GR}
R.A. Askey and J.A. Wilson, Mem.Amer.Math Soc 54 (319) (1985) \\
G. Gasper and M. Rahman, Basic hypergeometric series, Cambridge University Press, 1990\\
A.S. Zhedanov, Theor.Math.Fiz. 89 (1991) 190\\
T. Koornwinder, SIGMA 3, 063 (2007)\\
B. Aneva, M.Chaichian, P.P.Kulish, J.Phys.A41 (2008) 135201,  arXiv:0804.1623


\bibitem{orthopols} Chuan-Tsung Chan, A.Mironov, A.Morozov and A.Sleptsov,
arXiv:1712.03155

\bibitem{MMS} A.Mironov, A.Morozov and A.Sleptsov, Pis'ma v ZhETF, 106 (2017) 607, arXiv:1709.02290

\bibitem{arbor} A.Mironov, A.Morozov, An.Morozov, P.Ramadevi, V.K. Singh,
JHEP {\bf 1507} (2015) 109,  arXiv:1504.00371\\
A.Mironov and A.Morozov, Phys.Lett. B755 (2016) 47-57, arXiv:1511.09077 \\
A. Mironov, A. Morozov, An. Morozov, P. Ramadevi, V.K. Singh and A. Sleptsov,
J.Phys. A: Math.Theor. {\bf 50} (2017) 085201, arXiv:1601.04199

%\bibitem{RAS} Racah/AW pols for N=2???


\bibitem{M3141} A.Morozov,  Nucl.Phys. B911 (2016) 582-605,
arXiv:1605.09728

\bibitem{Mfact} A.Morozov,  JHEP 1609 (2016) 135,
arXiv:1606.06015 v8

\bibitem{KMsup} Ya.Kononov and A.Morozov,
Theor.Math.Phys. 193 (2017) 1630-1646,
arXiv:1609.00143

\bibitem{KMtwist} Ya.Kononov and A.Morozov,
Mod.Phys.Lett. A Vol. 31, No. 38 (2016) 1650223,
arXiv:1610.04778

\bibitem{M333} A.Morozov,  arXiv:1711.09277

\bibitem{Mnonrect} A.Morozov,
arXiv:1612.00422

\bibitem{MnonrectS} A.Morozov, Phys.Lett. B 766 (2017) 291-300,
arXiv:1701.00359

\bibitem{IMMMfe}
H. Itoyama, A. Mironov, A. Morozov and An. Morozov, JHEP 2012 (2012) 131, arXiv:1203.5978

\bibitem{evo} A. Mironov, A. Morozov and An. Morozov,
AIP Conf. Proc. 1562 (2013) 123, arXiv:1306.3197

\bibitem{diffexpan}
S.Arthamonov, A.Mironov, A.Morozov, Theor.Math.Phys. 179 (2014) 509-542, arXiv:1306.5682 \\
S.Arthamonov, A.Mironov, A.Morozov, An.Morozov, JHEP 04 (2014) 156, arXiv:1309.7984 \\
Ya.Kononov and A.Morozov, JETP Letters 101 (2015) 831-834,  arXiv:1504.07146\\
C.Bai, J.Jiang, J.Liang, A.Mironov, A.Morozov, An.Morozov, A.Sleptsov,
 arXiv:1709.09228

\bibitem{DMMSS}  P.Dunin-Barkowski, A.Mironov, A.Morozov, A.Sleptsov, A.Smirnov,
JHEP 03 (2013) 021, arXiv:1106.4305

\bibitem{MMMS22} A. Mironov, A. Morozov, An. Morozov, A. Sleptsov,
Phys.Lett. B760 (2016) 45-58,
arXiv:1605.04881

\end{thebibliography}

\documentclass{article}

\usepackage{amssymb,amsfonts,amsmath}
\usepackage{cite,enumerate,float,indentfirst}
\usepackage{color}

\def\be{\begin{eqnarray}}
\def\ee{\end{eqnarray}}
\def\nn{\nonumber}

\def\i{\mu}
\def\j{\nu}

\hoffset= - 1.0in         

\begin{document}

\hfill ITEP/TH-34/17

\hfill IITP/TH-21/17

\bigskip

\centerline{\Large{
Generalized hypergeometric series for Racah matrices in rectangular representations
}}

\bigskip

\bigskip

\centerline{\bf  A.Morozov }

\bigskip

{\footnotesize
\centerline{{\it
ITEP, Moscow 117218, Russia}}

\centerline{{\it
Institute for Information Transmission Problems, Moscow 127994, Russia
}}

\centerline{{\it
National Research Nuclear University MEPhI, Moscow 115409, Russia
}}
}

\bigskip

\bigskip

\centerline{ABSTRACT}

\bigskip

{\footnotesize
One of spectacular results in mathematical physics is the
expression of Racah matrices for symmetric representations of the quantum group $SU_q(2)$
through the Askey-Wilson polynomials,
associated with the $q$-hypergeometric functions ${_4\phi_3}$.
Recently it was shown that this is in fact the general property of symmetric representations,
valid for arbitrary $SU_q(N)$ -- at least for exclusive Racah matrices $\bar S$.
The natural question then is what substitutes the conventional $q$-hypergeometric polynomials
when representations are more general?
New advances in the theory of matrices $\bar S$, provided by the study of
differential expansions of knot polynomials, suggest that these are multiple sums
over Young sub-diagrams of the one, which describes the original representation of $SU_q(N)$.
A less trivial fact is that the entries of the sum are not just the factorized
combinations of quantum dimensions, as in the ordinary hypergeometric series,
but involve non-factorized quantities, like the skew characters and their further
generalizations -- 
as well as associated additional summations
with the Littlewood-Richardson weights.
}

\bigskip

\bigskip

\section{Introduction}

Racah matrices \cite{racahmatrices} describe the deviation from associativity in the product
of representations and they play a prominent and increasing role in modern
quantum field and string theory.
Despite in many cases we need these quantities for sophisticated representations
of infinite-dimensional algebras, they are still far from being well known even
for the simplest quantum algebras $SU_q(N)$.
In these simple cases Racah matrices are responsible for at least two
subjects of primary importance: modular transformations of conformal blocks \cite{modtrans}
in $2d$ conformal field theory \cite{CFT} and for calculation of knot invariants \cite{knotinvs}
(knot polynomials) in $3d$ Chern-Simons theory \cite{CS}.
Part of the problem is that Racah matrices are actually maps,
and explicit formulas are available in particular basises -- what often
makes these formulas non-invariant (this is often referred to as the
multiplicity problem), and thus not-very-interesting for pure mathematicians.
Thus the progress in the field is largely due to physical methods and
inspirations -- while the rigorous presentation awaits the completion
of the phenomenological part of the story.

In the present paper we open one more new chapter of this exciting mystery-book:
the relation between Racah matrices and hypergeometric orthogonal polynomials
of the Askey-Wilson type \cite{GR}.
Recently we reviewed the subject from the point of view of orthogonal polynomials
\cite{orthopols}, and now we approach it again from the Racah matrix side --
with the natural question: what generalizes hypergeometric series if we switch
from symmetric to generic representations of $SU_q(N)$.

As the first step
we demonstrated in \cite{MMS} that exclusive Racah matrices $\bar S_{\mu\nu}^{R}$,
$$
\Big((R\otimes \bar R)\otimes R \longrightarrow R \Big)
\ \stackrel{\bar S}{\longrightarrow} \
\Big(R\otimes (\bar R \otimes R) \longrightarrow R \Big)
$$
needed for arborescent knot calculus of \cite{arbor},
in the case of symmetric representations $R=[r]$
are expressed through the hypergeometric Racah/Askey-Wilson polynomials
for arbitrary algebra $SU_q(N)$ (for $N=2$ this is the classical result).

Not surprisingly, similar formulas follow from the general
$\bar S$-calculus of {\cite{M3141}--\!\!\!\cite{MnonrectS}}.
Since this calculus is applicable to arbitrary representations $R$,
and is especially well understood for rectangular $R=[r^s]$,
this opens a possibility to generalize the hypergeometric realization.
Relevant substitute of the hypergeometric series in this case are
sums over Young diagrams with the entries made from the skew characters of $su_q(N)$.

We begin from reminding the general ideas of $\bar S$-calculus in sec.\ref{dbfact}
and then describe in sec.\ref{Ffns1} what is currently known about the structure of the
underlying $F$-functions
and their relation \cite{KMtwist} to skew Schur polynomials.
After that in secs.\ref{symreps} and \ref{AW} we explain what happens in
symmetric representations $R=[r]$
and  how the Askey-Wilson realization arises for them in this context.
We conclude in sec.\ref{conc} with the suggestion that the formulas in sec.\ref{dbfact}
are exactly the ones, which provide the extension of Askey-Wilson realization
from symmetric to rectangular representations.
Non-rectangular case is also straightforward, but there are still some technical
difficulties to be resolved to handle non-trivial multiplicities -- the most
interesting part of both Racah and arborescent calculi.
We refer to \cite{Mnonrect,MnonrectS} for explanations and
leave non-rectangular case for the future considerations.

\section{Factorization of differential expansion for double braids
\label{dbfact}
}

According to \cite{M3141,Mfact} and \cite{KMsup,KMtwist} the rectangularly-colored
HOMFLY for a double braid $(m,n)$

\begin{picture}(200,280)(-230,-230)
\qbezier(-40,0)(-50,20)(-60,0)
\qbezier(-40,0)(-50,-20)(-60,0)
\qbezier(-20,0)(-30,20)(-40,0)
\qbezier(-20,0)(-30,-20)(-40,0)
\qbezier(-20,0)(-15,10)(-10,10)
\qbezier(-20,0)(-15,-10)(-10,-10)
\put(-5,0){\mbox{$\ldots$}}
\qbezier(10,10)(15,10)(20,0)
\qbezier(10,-10)(15,-10)(20,0)
\qbezier(20,0)(30,20)(40,0)
\qbezier(20,0)(30,-20)(40,0)
\qbezier(40,0)(50,20)(60,0)
\qbezier(40,0)(50,-20)(60,0)
\put(-60,0){\line(-1,2){10}}
\put(-60,0){\line(-1,-2){10}}
\put(60,0){\line(1,2){10}}
\put(60,0){\line(1,-2){10}}
\qbezier(0,-80)(-20,-90)(0,-100)
\qbezier(0,-80)(20,-90)(0,-100)
\qbezier(0,-100)(-20,-110)(0,-120)
\qbezier(0,-100)(20,-110)(0,-120)
\qbezier(0,-120)(-10,-125)(-10,-130)
\qbezier(0,-120)(10,-125)(10,-130)
\put(0,-145){\mbox{$\vdots$}}
\qbezier(0,-160)(-10,-155)(-10,-150)
\qbezier(0,-160)(10,-155)(10,-150)
\qbezier(0,-160)(-20,-170)(0,-180)
\qbezier(0,-160)(20,-170)(0,-180)
\qbezier(0,-180)(-20,-190)(0,-200)
\qbezier(0,-180)(20,-190)(0,-200)
\put(0,-80){\line(-2,1){10}}
\put(0,-80){\line(2,1){10}}
\put(0,-200){\line(-2,-1){10}}
\put(0,-200){\line(2,-1){10}}
\put(0,-200){\line(-2,-1){20}}
\put(0,-200){\line(2,-1){20}}
\qbezier(-10,-75)(-80,-40)(-70,-20)
\qbezier(10,-75)(80,-40)(70,-20)
\put(-10,-205){\vector(2,1){2}}
\put(10,-205){\vector(2,-1){2}}
\put(-65,10){\vector(-1,2){2}}
\put(65,10){\vector(-1,-2){2}}
\put(-70,-20){\vector(1,2){2}}
\put(70,-20){\vector(1,-2){2}}
\put(-3,20){\mbox{\footnotesize$2n$}}
\put(-32,-140){\mbox{\footnotesize $2m$}}
%
\qbezier(-70,20)(-80,40)(-97,25)
\qbezier(-97,25)(-111,13))(-100,-30)
\qbezier(-100,-30)(-60,-230)(-20,-210)
\qbezier(70,20)(80,40)(97,25)
\qbezier(97,25)(111,13))(100,-30)
\qbezier(100,-30)(60,-230)(20,-210)
\put(-102,-22){\vector(1,-4){2}}
\put(100,-30){\vector(1,4){2}}
\end{picture}

\noindent
is given by a {\it factorized} differential expansion {\cite{IMMMfe}-\!\!\cite{diffexpan}}:
\be
H_R^{(m,n)}
=  \sum_{\mu,\nu\subset R}
\frac{\sqrt{{\cal D}_\mu{\cal D}_\nu}}{d_R}\,\bar S_{\mu\nu}^{R}\, \Lambda_\mu^m\Lambda_\nu^n
=  \sum_{\lambda\subset R} \chi^*_{\lambda^{tr}}(r)\chi^*_\lambda(s)\cdot
\{q\}^{2|\lambda|} h_\lambda^2\, \chi^*_\lambda(N+r)\chi^*_\lambda(N-s) \cdot
\frac{F_\lambda^{(m)} F_\lambda^{(n)}}{F_\lambda^{(1)}F_\lambda^{(-1)}}
\label{HRdb}
\ee
Here $\{x\} = x-x^{-1}$, quantum numbers are $[n]=\frac{\{q^n\}}{\{q\}}$ and
$h_\lambda$ is denominator in
\be
\chi^*_\lambda(N) =
{\rm Schur}\!\left\{p_k=\frac{[Nk]}{[k]}\right\} =
\frac{1}{h_\lambda} \prod_{(\alpha,\beta)\in \lambda}
[N+\alpha-\beta]
\ee
given by the hook formula:
\be
h_\lambda = \prod_{(\alpha,\beta)\in\lambda} [{\rm hook\ length}_{(\alpha,\beta)}]
= \prod_{(\alpha,\beta)\in\lambda} [{\rm leg}_{(\alpha,\beta)} + {\rm arm}_{(\alpha,\beta)}+1]
\ee
For the figure-eight knot $4_1$ parameters are $(m,n)=(1,-1)$,
and the last factor (the ratio of four $F$) is absent.
Generic twist knots correspond to $n=1$, and, since actually $F_\lambda^{(-1)}=1$
the last ratio turns into just $F^{(m)}_\lambda$.

The difference between knot and Racah calculi
is that for the latter one we need not the $F^{(m)}$ as a total for a given $m$,
but its decomposition into particular eigenvalues, dictated by the evolution method
of \cite{DMMSS} and \cite{evo}.
Among other things this eliminates the factor $\{q\}^{|\lambda|}h_\lambda$ from
the expression for $\bar S$,
making it from just the characters $\chi^*$ at the topological locus \cite{DMMSS}
(quantum dimensions) with a small admixture of additional quantum-number factors.

$F_\lambda^{(m)}$ are actually sums over $m$-the powers of the squared "eigenvalues"
$\Lambda_\mu^{m}$, where $\mu$ are Young sub-diagrams of $\lambda$ --
and according to \cite{Mfact} one can read from this eigenvalue expansion the Racah matrices
$\bar S^{R}_{\mu\nu}$.
According to \cite{KMtwist} the coefficients of the $F$ expansion are essentially
the skew characters $\chi_{\lambda/\mu}=\sum_\nu C^\lambda_{\mu\nu}\chi_\nu$
with the Littlewood-Richardson coefficients $C^\lambda_{\mu\nu}$, defined from
$\chi_\mu\chi_\nu = \sum_\lambda C^\lambda_{\mu\nu} \chi_\lambda$.

Eq.(\ref{HRdb}) and the general formula for $F$-expansion in \cite{KMtwist}
implies  for the case of rectangular representation $R=[r^s]$ with $s\leq 2$ that
\be
\!\!\!
\boxed{
\frac{\sqrt{{\cal D}_\mu{\cal D}_\nu}}{d_R}\,\bar S_{\mu\nu}^{[r^s]}\!
= \!\!\!\!\!\!\!\sum_{\mu,\nu\subset \lambda\subset R}\!\!\!\!\!\!\!\!
\,\frac{(-)^{|\lambda|}
\chi^*_{\lambda^{tr}}(r) \chi^*_\lambda(s) \,\chi^*_\lambda(N+r)\chi^*_\lambda(N-s)}
{\chi^*_\lambda(N)^2}
\!\cdot\! G^\mu_\lambda G^\nu_\lambda\!\! \cdot\!
\frac{\!\chi_{\lambda/\mu}^*(N_\mu^\lambda)\,\chi_{\mu^{tr}}^*(N_\mu^\lambda)}
{\chi_\lambda^*(N_\mu^\lambda)}
\!\cdot\! \frac{\!\chi_{\lambda/\nu}^*(N_\nu^\lambda)\,\chi_{\nu^{tr}}^*(N_\nu^\lambda)}
{\chi_\lambda^*(N_\nu^\lambda)}
}\!\!\!\!
\label{bSexpansion}
\ee
where
$N_\mu^\lambda = N + \sum i - \sum j + \lambda\text{-dependent shift}\ $
for the hook-parametrization \cite{Mfact} of $\mu = (i_1,j_1|i_2,j_2|\ldots )$
and $G_\lambda^\mu$ are some factorized quantities:
ratios of quantum numbers $\ [\,N+u\,]\ $ with various shifts $u$,
also described in terms of the hook parameters $\{i,j\}$ for $\mu$ and $\{a,b\}$ for $\lambda$.
As to ${\cal D}_\mu$ at the l.h.s., they are dimensions of representations,
appearing in decomposition of the product $R\otimes \bar R$ --
which for rectangular $R$ are in one-to-one correspondence with the Young sub-diagrams
of $R$ \cite{KMsup}.

\section{$F$-functions and $G$-factors
\label{Ffns1}}

According to \cite{Mfact} and \cite{KMtwist} the $F$-functions are best described
in a peculiar hook parametrization of Young diagrams:

\begin{picture}(300,150)(-50,-10)
\put(0,0){\line(0,1){130}}
\put(15,0){\line(0,1){130}}
\put(30,15){\line(0,1){85}}
\put(45,30){\line(0,1){40}}
\put(75,30){\line(0,1){15}}
\put(90,15){\line(0,1){15}}
\put(120,0){\line(0,1){15}}

\put(0,0){\line(1,0){120}}
\put(0,15){\line(1,0){120}}
\put(15,30){\line(1,0){75}}
\put(30,45){\line(1,0){45}}
\put(0,130){\line(1,0){15}}
\put(15,100){\line(1,0){15}}
\put(30,70){\line(1,0){15}}

\put(3,50){\mbox{$a_1$}}
\put(18,55){\mbox{$a_2$}}
\put(33,60){\mbox{$a_3$}}
\put(50,3){\mbox{$b_1$}}
\put(55,18){\mbox{$b_2$}}
\put(60,33){\mbox{$b_3$}}

\put(130,100){\mbox{a 3-hook Young diagram $\ (a_1,b_1|a_2,b_2|a_3,b_3)$}}
\put(130,80){\mbox{$= \, [a_1+1,a_2+2,a_3+3,3^{b_3},2^{b_2-b_3-1},1^{b_1-b_2-1}]$}}
\put(130,60){\mbox{of the size $\ a_1+a_2+a_3+b_1+b_2+b_3+3$}}

\end{picture}

\noindent
The main drawback of this parametrization is that it changes discontinuously
with the number of hooks:
the empty diagram $\emptyset$ is not a particular case of any 1-hook diagram
$(a_1,b_1)$, of which the minimal is $[1]=(0,0)$ and so on.
Formally one could associate $\emptyset$ with $a_1+b_1=-1$, but this is not
quite respected by the formulas.
Because of this one needs to write
\be
F_\lambda^{(m)}(A,q) = \frac{c_\lambda }{\{q\}^{|\lambda|}h_\lambda\cdot \chi^*_\lambda(N)  }
\sum_{\mu\subset\lambda} f_\lambda^\mu(N,q) \cdot \Lambda_\mu^m
\label{Fvsf}
\ee
with different expressions for different
hook numbers $\#^h_\lambda$ and $\#^h_\mu$.
The $F$-functions depend explicitly on $A=q^N$ and $q$, but mostly are made
from the quantum numbers, involving $N$.
The only exceptions are the squared eigenvalues
\be
\Lambda_\mu = \Lambda_{(i_1,j_1|i_2,j_2|\ldots)} =
\prod_{k=1}^{\#^h_\mu} (A\cdot q^{i_k-j_k})^{2(i_k+j_k+1)}
\ee
and the overall coefficients
\be
c_\lambda = c_{(a_1,b_1|a_2,b_2|\ldots)} =
\prod_{k=1}^{\#^h_\lambda} (A\cdot q^{\frac{a_k-b_k}{2}})^{(a_k+b_k+1)}
\ee
Both, however, drop away from the expression (\ref{HRdb}) for the
Racah matrix $\bar S^R $ -- $\Lambda_\mu$ because $\bar S^R_{\mu\nu}$
are coefficients of the $\Lambda$-expansion and $c_\lambda$ because of
cancellations, dictated by the properties:
\be
F_\lambda^{(-1)}=1, \ \ \ \ \ F_\lambda^{(0)}=\delta_{\lambda,\emptyset}, \ \ \ \ \
F_\lambda^{(1)} = (-)^{\sum_{k}  (a_k+b_k+1)}c_\lambda^2
\label{sumrules}
\ee
which are responsible for the simplicity of the differential expansion
{\cite{IMMMfe}-\!\!\cite{diffexpan}}
at the r.h.s. of  (\ref{HRdb})
for respectively the figure-eight knot $4_1$, unknot and   the trefoil $3_1$.
As already mentioned after (\ref{HRdb}), the factors $\{q\}$ and $h_\lambda$
also drop away from the expressions for HOMFLY polynomials and $\bar S$.

The sum rules (\ref{sumrules}) are non-trivial analogues of the elementary identity
\be
\sum_{\mu\subset \lambda} (-)^{|\mu|}\cdot\chi_{\lambda/\mu}\cdot \chi_{\mu^{tr}}
= \delta_{\lambda,\emptyset}
\label{naivesumrule}
\ee
which follows from the defining property of skew characters,
\be
\sum_{\mu\subset \lambda} \chi_{\lambda/\mu}\{p'_k\}\cdot \chi_{\mu^{tr}}\{p''_k\} =
\chi_\lambda\{p'_k+p''_k\}
\ee
\vspace{-0.6cm}

\noindent
and the transposition law
\vspace{-0.4cm}
\be
\chi_\mu\{-p_k\} = (-)^{|\mu|} \chi_{\mu^{tr}}\{p_k\}
\ee
While (\ref{naivesumrule}) holds beyond the topological locus (i.e. for all values
of time variables), it does not survive introduction of weights $\Lambda^{\pm 1}$
even on the locus,
i.e. there is no analogue of the other two identities in (\ref{sumrules}).

The difficult part of the story is to describe $f_\lambda^\mu$ which satisfy all the three.
Currently they are fully known for $\lambda=(a_1,b_1|a_2,0)$ --
what is enough to get the Racah matrices $\bar S$ for the case $R=[r,r]$
(actually, for this purpose $b_1=0,1$ is sufficient).
After (\ref{naivesumrule}) it is not such a big surprise that they involve skew characters,
but exact formulas \cite{Mfact,KMtwist} are still not very easy to interpret and understand.

$\bullet$ For the empty diagram $\mu$ always
\vspace{-0.3cm}
\be
f_\lambda^\emptyset = 1
\ee

$\bullet$ Since $\mu\subset\lambda$ the number of hooks $\#^h_\mu\leq \#^h_\lambda$.
Thus for the single-hook $\lambda$ it remains to describe only the
contributions of the single-hook $\mu$.
These are relatively simple factorized expressions \cite{Mfact}:
\be
f_{(a,b)}^{(i,j)} = g_{(a,b)}^{(i,j)} \cdot K_{(a,b)}^{(i,j)}
= (-)^{i+j+1}
\cdot \frac{[a]!}{[a-i]![i]!}\cdot\frac{[b]!}{[b-j]![j]!}\cdot\frac{[a+b+1]}{[i+j+1]}
\cdot \frac{D_a!D_i!}{D_{a+i+1}!}\cdot\frac{\bar D_b!\bar D_j!}{\bar D_{b+j+1}!}
\cdot \frac{D_{2i+1}D_{-2j-1}}{D_0D_{i-j}}
\label{f1vsgK}
\ee
with
\vspace{-0.5cm}
\be
g_{(a,b)}^{(i,j)}
= (-)^{i+j+1} \cdot \frac{D_{2i+1}\bar D_{2j+1}}{D_0D_{i-j}}\cdot
\frac{(D_a!)^2}{D_{a+i+1}!\,D_{a-i-1}!}\cdot\frac{(\bar D_b!)^2}{\bar D_{b+j+1}!\,\bar D_{b-j-1}!}
\label{gfactors}
\ee
and
\vspace{-0.5cm}
\be
K_{\lambda}^{\mu}(N) = \frac{\chi^*_{\lambda/\mu}(N)\,\chi^*_{\mu}(N)}{\chi^*_{\lambda}(N)}
\ee
Note that this combination involves $\chi_\mu$ rather than $\chi_{\mu^{tr}}$,
thus $\sum_\mu (-)^{|\mu|}K_{\lambda}^\mu \neq 0$ (in fact, it vanishes, but only for
diagrams $\lambda$ of odd size $|\lambda|=odd$, because
$(-)^{|\mu|}\chi_\mu\{p_k\} = \chi_\mu\{(-)^kp_k\} $, i.e. only odd times change sign).
Notation in (\ref{gfactors}) is: $D_a=[N+a]$, $\bar D_b=[N-b]$ and $D_a!=\prod_{k=0}^a D_k
= \frac{[N+a]!}{[N-1]!}$,
$\bar D_b! = \prod_{k=0}^b \bar D_k = \frac{[N]!}{[N-b-1]!}$
(note that these products start from $k=0$ and include
respectively $a+1$ and $b+1$ factors).

$\bullet$ For two-hook $\lambda=(a_1,b_1|a_2,b_2)$ the formulas are far more involved,
and they are different for different number of hooks in $\mu$:
\be
f_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1)}
= f^{(a_1,b_1)}_{(i_1,j_1)} \cdot \xi_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1)}
= g_{(a_1,b_1)}^{(i_1,j_1)}\cdot K_{(a_1,b_1)}^{(i_1,j_1)}(N)
\cdot \xi_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1)}
\ee
\be
f_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1|i_2,j_2)} =
\frac{[N+i_1+i_2+1][N-j_1-j_2-1]}{[N+i_1-j_2][N+i_2-j_1]}
\cdot \underbrace{g_{(a_1,b_1)}^{(i_1,j_1)}\,g_{(a_2,b_2)}^{(i_2,j_2)}
\cdot K_{(a_1,b_1 )}^{(i_1,j_1 )}(N)
K_{( a_2,b_2)}^{( i_2,j_2)}(N)}_{f_{(a_1,b_1)}^{(i_1,j_1)}\,f_{(a_2,b_2)}^{(i_2,j_2)}}
\cdot\, \xi_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1|i_2,j_2)}
\ee
Non-trivial are the correction factors:
\be
\xi_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1)} =
\  \boxed{\delta_{a_2\cdot b_2,0}}\cdot
\left(
\frac{[N+a_2-j_1][N-b_2+i_1]}{[N+a_2+i_1+1][N-b_2-j_1-1]}
\cdot  {\frac{K_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1)}(N+i_1-j_1) }
{K_{(a_1,b_1)}^{(i_1,j_1)}(N+i_1-j_1)}}
\cdot \boxed{\delta_{i_1\cdot j_1,0}}\ +
\right.
\nn\ee
\vspace{-0.4cm}
\be
\!\!\!\!\!
+ \
\left.
 { \frac{K_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1)}
 \Big(N+(i_1+1)\delta_{b_2,0}-(j_1+1)\delta_{a_2,0}\Big)
 }
{K_{(a_1,b_1 )}^{(i_1,j_1)}\Big(N+(i_1+1)\delta_{b_2,0}-(j_1+1)\delta_{a_2,0}\Big)}
}
\cdot\ \boxed{(1-\delta_{i_1,0})(1-\delta_{j_1,0})}
\right)
+ \ \boxed{(1-\delta_{a_2,0})(1-\delta_{b_2,0}) \ \cdot \, ?}
\label{xi1}
\ee
and
\be
\!\!\!
\xi_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1|i_2,j_2)} =
{\frac{\boxed{\delta_{a_2\cdot b_2,0}}\cdot K_{(a_1,b_1|a_2,b_2)}^{(i_1,j_1|i_2,j_2)}
\Big(N+(i_1+i_2+2)\cdot \delta_{b_2,0} - (j_1+j_2+2)\cdot \delta_{a_2,0}\Big)
}
 {\Big( K_{(a_1,b_1 )}^{(i_1,j_1)}
 \cdot K_{(a_2,b_2 )}^{(i_2,j_2)}\Big)
 \Big(N+(i_1+i_2+2)\cdot \delta_{b_2,0} - (j_1+j_2+2)\cdot \delta_{a_2,0}\Big)}}
 + \ \boxed{(1-\delta_{a_2,0})(1-\delta_{b_2,0}) \ \cdot \, ?}
\nn\ee
Note that we provide expressions only for the case when $a_2\cdot b_2=0$
(i.e. when either $b_2=0$ or $a_2=0$),
what is emphasized by boxes in above formulas.
Sufficient for all the simplest non-symmetric rectangular representations
$R=[r,r]$ and $R=[2^r]$ are respectively $b_2=0$ and $a_2=0$.
Note also, that $K$ factorizes nicely for the single-hook diagrams $\lambda$:
\be
K_{(a,b)}^{(i,j)}(N) =
\frac{\chi^*_{(a,b)/(i,j)}(N)\cdot \chi^*_{(i,j) }(N)}{\chi^*_{(a,b)}(N)} =\nn \\
= \frac{[a]!}{[i]!\,[a-i]!}\cdot\frac{[b]!}{[j]!\,[b-j]!}\cdot \frac{[a+b+1]}{[i+j+1]}
\cdot[N]\cdot \frac{[N+i]!}{[N-j-1]!}\cdot\frac{[N+a-i-1]!}{[N+a]!}
\frac{ [N-b-1]!}{[N-b+j]!}
\label{K1}
\ee
but does not do so
for the two-hook $\lambda$, even if $\mu$ is still a single-hook,
like in (\ref{xi1}), e.g.
\be
K_{(3,1|1,0)}^{(1,1)}(N) \sim  \chi^*_{[43]/[21]}(N)
= \chi^*_{[31]}(N)+\chi^*_{[22]}(N) \sim \nn \\
\sim
A^2q^8+2A^2q^6+A^2q^4+A^2q^2-q^6-q^4-2q^2-1 \sim
[3][2][N+2]+[4][N] \sim
[2]^2[N+2]+[N-2]
\ee
Therefore there seems to be no freedom to change the somewhat mysterious
shifts of $N$ in these formulas.

In the case of $a_2\cdot b_2\neq 0$ skew characters are further deformed
into still more complicated quantities \cite{M333}, which still lack a proper
identification.

\section{Symmetric representations
\label{symreps}}

For symmetric representations $R=[r]$ with $s=1$
we obtain from (\ref{HRdb}), by recursively substituting
(\ref{Fvsf}), (\ref{f1vsgK}), (\ref{gfactors}) and (\ref{K1}):
\be
\bar\sigma_{\mu\nu}^{[r]} =  d_r\cdot
 \frac{[\mu]! [\nu]![N-1]![N-2]!}{[N+\mu-2]![N+\nu-2]!}\,
\cdot\!\!\!\!\sum_{\mu,\nu\leq \lambda\leq r}\!\!\!\!(-)^\lambda
\frac{[\lambda]!}{[r-\lambda]!\,[\lambda-\mu]!\,[\lambda-\nu]!}
\cdot \frac{[N+r+\lambda-1]![N+\lambda-2]!}{[N+\lambda+\mu-1]![N+\lambda+\nu-1]!}
= \nn
\ee
\vspace{-0.5cm}
\be
= \frac{[r]!}{[\mu]![\nu]!}\,
\frac{\chi_r^*(N)^2}{\chi^*_\mu(N)\chi^*_\mu(N-1)\,\chi^*_\nu(N)\chi^*_\nu(N-1)}\cdot
\sum_{\lambda={\rm max}(\mu,\nu)}^{r}
\frac{(-)^\lambda\cdot [\lambda]!}{[r-\lambda]![\lambda-\mu]![\lambda-\nu]!}
\,\frac{ \chi^*_\lambda(N+r)\chi^*_\lambda(N-1)}{\chi^*_\lambda(N+\mu)\chi^*_\lambda(N+\nu)}
\label{symrepser}
\ee
where used are also explicit expression for dimensions
\be
{\cal D}_\mu =  {[N+2\mu-1][N-1]}
\left(\frac{[N+\mu-2]!}{[\mu]!\,[N-1]!}\right)^2
\ee
of representations in
$R\otimes \bar R = [r]\otimes [r^{N-1}] =
\oplus_{\mu=0}^r [r+\mu,r^{N-2},r-\mu]=\oplus_{\mu=0}^r [2\mu,\mu^{N-2}]$
to convert the original $\ \frac{\sqrt{{\cal D}_\mu{\cal D}_\nu}}{d_r}\,\bar S_{\mu\nu}^{[r]}\ $
into $\ \bar\sigma_{\mu\nu}^{[r]} = (-)^{\mu+\nu} \frac{d_r}{\sqrt{{\cal D}_\mu{\cal D}_\nu}}
\,\bar S_{\mu\nu}^{[r]}\ $.

\bigskip

This expression is partly in terms of quantum dimensions $\chi^*_\lambda$ and does contains
neither skew characters, nor shifts.
The only thing which reminds that the skew characters are somewhere behind the scene
is the presence of factorials $\ [\lambda-\mu]!\,[\lambda-\nu]!\ $ in denominator.
Of course, it is possible to make this explicit, by
rewriting (\ref{symrepser}) in the form (\ref{bSexpansion}):
\be
\!\!\!
\bar\sigma_{\mu\nu}^{[r]} =
\frac{1}{\chi^*_\i(N-1)\,\chi^*_\j(N-1)}\ \cdot \!\!\!\!\!\!
\sum_{\lambda={\rm max}(\mu,\nu)}^{r}
\boxed{(-)^\lambda \chi_r^*(N)\chi^*_r(N+\lambda)\cdot
\frac{ [r]!}{[\lambda]![r-\lambda]!} \cdot\frac{[N-1]}{[N+\lambda-1]}}
\cdot\!
\frac{\chi^*_{\lambda-\mu}(N+\mu)\,\chi^*_{\lambda-\nu}(N+\nu)}
{\chi^*_\lambda(N+\mu)\,\chi^*_\lambda(N+\nu)}
\nn
\ee
with relatively simple  shifts
$N_{[\mu]}^{[\lambda]} = N+\mu-1$, which are independent of $\lambda$,
and  $G$-factors
$G_{[\lambda]}^{[\mu]}=(-)^{\mu}\frac{[N+\lambda-1][N+2\mu-1]}{[N+\lambda+\mu-1][N+\mu-1]}$.

\bigskip

{\bf Example of $R=[1]$:}

\bigskip


In we denote the combination in the box in the last formula
through $B_\lambda$, then

\be
\bar\sigma^{[1]}=  \left(\begin{array}{ccc}
\frac{1}{\chi_0(N-1)\chi_0(N-1)}\left(\frac{B_0 \chi_0(N)\chi_0(N)}{\chi_0(N)\chi_0(N)}
+ \frac{B_1 \chi_1(N)\chi_1(N)}{\chi_1(N)\chi_1(N)} \right) &&
\frac{1}{\chi_0(N-1)\chi_1(N-1)}\cdot \frac{B_1 \chi_1(N)\chi_0(N+1)}{\chi_1(N)\chi_1(N+1)}\\ \\
\frac{1}{\chi_1(N-1)\chi_0(N-1)}\cdot \frac{B_1 \chi_0(N+1)\chi_1(N)}{\chi_1(N+1)\chi_1(N)} &&
\frac{1}{\chi_1(N-1)\chi_1(N-1)}\cdot \frac{B_1 \chi_0(N+1)\chi_0(N+1)}{\chi_1(N+1)\chi_1(N+1)}
\end{array}\right) = \nn \\ \nn \\ \nn \\
\!\!\!\!\!\!\!\!\!\!\! =  \left(\begin{array}{ccc}
  B_0+ B_1   &&
\frac{1}{ \chi_1(N-1)}\cdot \frac{B_1  }{ \chi_1(N+1)}\\ \\
\frac{1}{\chi_1(N-1) }\cdot \frac{B_1   }{\chi_1(N+1) } &&
\frac{1}{\chi_1(N-1)\chi_1(N-1)}\cdot \frac{B_1  }{\chi_1(N+1)\chi_1(N+1)}
\end{array} \right) =
\left(\begin{array}{ccc}  1 && -1 \\ \\ -1 &&
-\frac{1}{[N+1] [N-1]}\end{array}\right)
\ \ \ \ \
\ee
with $B_1=- {[N+1][N-1]} $ and $B_0 =  {[N]^2}$.

After multiplication by $\frac{(-)^{i+j}\sqrt{{\cal D}_i{\cal D}_j}}{d_{[r]}}$
with ${\cal D}_0=1$,
${\cal D}_1 = [N+1][N-1]$  and $d_{[1]}=[N]$ this gives the unitary symmetric Racah matrix
\be
\bar S_{[1]} = \frac{1}{[N]} \left(\begin{array}{ccc}
1 & & \sqrt{[N+1][N-1]} \\ \\ \sqrt{[N+1][N-1]} && -1 \end{array}\right)
\ee

\bigskip

{\bf Example of $R=[2]$:}

\bigskip


Similarly, in this case

\bigskip

\centerline{
{\tiny
$
\left(\!\!\!\!\begin{array}{ccc}
\frac{1}{\chi_0(N-1)\chi_0(N-1)}\left(\frac{B_0 \chi_0(N)\chi_0(N)}{\chi_0(N)\chi_0(N)}
+ \frac{B_1 \chi_1(N)\chi_1(N)}{\chi_1(N)\chi_1(N)}
+\frac{B_2 \chi_2(N)\chi_2(N)}{\chi_2(N)\chi_2(N)}\right) &
\frac{1}{\chi_0(N-1)\chi_1(N-1)}\left( \frac{B_1 \chi_1(N)\chi_0(N+1)}{\chi_1(N)\chi_1(N+1)}
+ \frac{B_2 \chi_2(N)\chi_1(N+1)}{\chi_2(N)\chi_2(N+1)}\right)
& \!\!\!\!\!\!
\frac{1}{\chi_0(N-1)\chi_2(N-1)}\cdot \frac{B_2 \chi_2(N)\chi_0(N+2)}{\chi_2(N)\chi_2(N+2)}
\\ \\
\frac{1}{\chi_1(N-1)\chi_0(N-1)}\left(
\frac{B_1 \chi_0(N+1)\chi_1(N)   }{\chi_1(N+1)\chi_1(N)}+
\frac{B_2 \chi_1(N+1)\chi_2(N)}{\chi_2(N+2)\chi_2(N)}\right) &
\!\!\!\!\!\!\!\!\!\!\! \frac{1}{\chi_1(N-1)\chi_1(N-1)}\left(
\frac{B_1 \chi_0(N+1)\chi_0(N+1)   }{\chi_1(N+1)\chi_1(N+1)}+
\frac{B_2 \chi_1(N+1)\chi_1(N)}{\chi_2(N+2)\chi_2(N+1)}\right)&
\!\!\!\! \frac{1}{\chi_1(N-1)\chi_2(N-1)}\cdot
\frac{B_2 \chi_1(N+1)\chi_0(N+2)}{\chi_2(N+2)\chi_2(N+2)}
\\ \\
\frac{1}{\chi_2(N-1)\chi_0(N-1)}\cdot \frac{B_2 \chi_0(N+2)\chi_2(N)}{\chi_2(N+2)\chi_2(N)} &
\frac{1}{\chi_2(N-1)\chi_1(N-1)}\cdot \frac{B_2 \chi_0(N+2)\chi_1(N+1)}{\chi_2(N+2)\chi_2(N+1)}
& \!\!\!\!
\frac{1}{\chi_2(N-1)\chi_2(N-1)}\cdot \frac{B_2 \chi_0(N+2)\chi_0(N+2)}{\chi_2(N+2)\chi_2(N+2)}
\end{array}\!\!\!\!\right) = \!\!
$
}
}

\bigskip

\bigskip

\centerline{
{\footnotesize
$
=  \left(\begin{array}{ccc}
B_0+B_1+B_2 &
\frac{1}{\chi_1(N-1)}\left( \frac{B_1 }{\chi_1(N+1)}
+ \frac{B_2  \chi_1(N+1)}{ \chi_2(N+1)}\right)
&
\frac{1}{\chi_2(N-1)}\cdot \frac{B_2 }{\chi_2(N+2)}
\\ \\
\frac{1}{\chi_1(N-1)}\left(
\frac{B_1    }{\chi_1(N+1)}+
\frac{B_2 \chi_1(N+1)}{\chi_2(N+2)}\right) &
\frac{1}{\chi_1(N-1)\chi_1(N-1)}\left(
\frac{B_1    }{\chi_1(N+1)\chi_1(N+1)}+
\frac{B_2 \chi_1(N+1)\chi_1(N)}{\chi_2(N+2)\chi_2(N+1)}\right)&
\frac{1}{\chi_1(N-1)\chi_2(N-1)}\cdot
\frac{B_2 \chi_1(N+1)}{\chi_2(N+2)\chi_2(N+2)}
\\ \\
\frac{1}{\chi_2(N-1)}\cdot \frac{B_2 }{\chi_2(N+2)} &
\frac{1}{\chi_2(N-1)\chi_1(N-1)}\cdot \frac{B_2 \chi_1(N+1)}{\chi_2(N+2)\chi_2(N+1)}&
\frac{1}{\chi_2(N-1)\chi_2(N-1)}\cdot \frac{B_2 }{\chi_2(N+2)\chi_2(N+2)}
\end{array}\right) =
$
}
}

\bigskip


\be
=  \left(\begin{array}{ccc}
1 & -1 & 1 \\
-1 & \frac{[N+2][N]-[2]^2}{[2][N+2][N-1]}& \frac{[2]}{[N-1][N+2] } \\
1 &\frac{[2]}{[N-1][N+2] } & \frac{[2]^2}{[N-1][N][N+2][N+3]}
\end{array}\right)
\ee

\bigskip

\noindent
with $B_0=\frac{[N]^2[N+1]^2}{[2]^2}$, $B_1 = -\frac{[N+2][N+1]^2[N-1]}{[2]}$ and
$B_2 = \frac{[N-1][N][N+2][N+3]}{[2]^2}$.

\bigskip

Multiplication by $\frac{(-)^{i+j}\sqrt{{\cal D}_i{\cal D}_j}}{d_{[r]}}$
with ${\cal D}_0=1$,
${\cal D}_1 = [N+1][N-1]$, ${\cal D}_2 = \frac{[N+3][N]^2[N-1]}{[2]^2}$
and $d_{[2]}=\frac{[N][N+1]}{[2]}$
provides the unitary symmetric Racah matrix
\be
\bar S_{[2]} = \frac{[2]}{[N+1][N]}\left(\begin{array}{ccc}
1 & \sqrt{[N+1][N-1]} & \frac{[N]\sqrt{[N+3][N-1]}}{[2]} \\
\sqrt{[N+1][N-1]} & [N+1]\frac{[N+2][N]-[2]^2}{[2][N+2]}&
-\frac{[N]}{[N+2]}\sqrt{ {[N+3][N+1]}} \\
\frac{[N]\sqrt{[N+3][N-1]}}{[2]}
&-\frac{[N]}{[N+2]}\sqrt{ {[N+3]}{[N+1]}} &
\frac{[N]}{[N+2]}
\end{array}\right)
\ee

\section{Hypergeometric series
\label{AW}}

Coming back to (\ref{symrepser}),  one can make
a change of summation variable $\lambda =r-k$
to get a factorial $[k]!$ in the denominator.
Then the sum turns into
\be
\bar\sigma_{\mu\nu}^{[r]} =
\frac{[N-1]\,\chi^*_r(N)}{\chi^*_\mu(N-1)\chi^*_\nu(N-1)}
\!\!\!\!\!\!
\sum_{k=0}^{{\rm min}(r-\mu,r-\nu)}
\!\!\!\!\!\!
 \frac{(-)^{r-k} \,[r-k]!}{[k]![r-\mu-k]![r-\nu-k]!}
\frac{[N+2r-1-k]![N+r-2-k]!}{[N+r+\mu-1-k]![N+r+\nu-1-k]!}
\label{symrepser1}
\ee
what is proportional to the $q$-hypergeometric polynomial
\be
 {_4\phi_3} \left(\begin{array}{c}
\mu-r,\ \ \nu-r,\ \ 1-N-r-\mu,\ \ 1-N-r-\nu \\ -r,\ \ 1-2r-N,\ \ 2-r-N
\end{array}\Big|\, z=q^2\right)
\ee
This looks different but is actually equivalent to the result of \cite{MMS}
for the same $\bar\sigma_{\mu\nu}^{[r]}$:
$$
\!\!\!\!
\frac{([\i]![\j]!)^2[r-\i]![r-\j]!\,[N-1]![N-2]!}{[r+\i+N-1]![r+\j+N-1]!}\cdot
d_r \cdot \!\!\!\!\!\!\!
\sum_{k={\rm max}(\i,\j)}^{{\rm min}(r,\i+\j)}\!\!\!\!\!\!\! (-)^k\,
\frac{[k+N+r-1]!}{([k-\i]![k-\j]!)^2[r-k]![\i+\j-k]![\i+\j+N-k-2]!} \sim
$$
\vspace{-0.4cm}
\be
\sim {_4\phi_3} \left(\begin{array}{c}
\i-r,\ \ \i-r,\ \ \j-r,\ \ \j-r \\ 1-2r-N,\ \ \i+\j+1-r,\ \ \i+\j+N-r-1
\end{array}\Big|\,z=q^2\right)
\label{MMSfla}
\ee
Note that (\ref{symrepser}) and (\ref{symrepser1}) contain just five factorials in denominator,
instead of  seven in the first line of (\ref{MMSfla}), which are usual
in the standard formulas for $SU_q(2)$ Racah matrices.

We remind \cite{orthopols} that in the balanced case, i.e. for $\alpha_1+\ldots+\alpha_{p+1}+1 =
\beta_1+\ldots+\beta_p$, hypergeometric series are expressed through quantum numbers and
\be
\!\!\!\!\!\!\!\!\!
{_{p+1}\phi_p}\left(\begin{array}{c}
\alpha_1\ \ldots\ \alpha_{p+1}\\ \beta_1\ \ldots\ \beta_p\end{array}\Big|\,z=q^2\right)
= \sum_n \frac{\chi^*_n(\alpha_1)\ldots \chi^*_n(\alpha_{p+1})}
{\chi^*_n(\beta_1)\ldots\chi^*_n(\beta_p)} = \nn \\
= \sum_n \frac{[\alpha_1+n-1]!\ldots [\alpha_{p+1}+n-1]!\,[\beta_1-1]!\ldots[\beta_p-1]!}
{[\alpha_1-1]!\ldots [\alpha_{p+1}-1]!\,[\beta_1+n-1]!\ldots [\beta_p+n-1]!\,[n]!}
\ \sim \ \sum_k \frac{[-\beta_1-n]!\ldots[-\beta_p-k]!}
{[-\alpha_1-k]!\ldots [-\alpha_{p+1}-k]!\,[k]!}
\ee

\section{Conclusion
\label{conc}}

In this sense (\ref{bSexpansion}) can be considered as the generalization of the
q-hypergeometric polynomials, which is relevant for description of generic Racah matrices,
at least in rectangular representations.
In non-rectangular case some sub-diagrams $\lambda\in R$ appear with non-trivial
multiplicities and contribute additional terms into this expansion,
see \cite{Mnonrect,MnonrectS} for more details.

In general the elements of Racah matrix $\bar S_{\mu\nu}^R$ are expressed through
quantum numbers, but are not factorized -- for two reasons:
because of the sum over sub-diagrams $\lambda\in R$ and because the items in the sum
are made not just from the nicely-factorized quantum dimensions
$\chi^*_\lambda, \chi^*_\mu, \chi^*_\nu$, but also
from the {\it skew} characters $\chi^*_{\lambda/\mu}$
and $\chi^*_{\lambda/\nu}$
and their further generalizations, like (23) and (31) in \cite{M333},
which do {\it not} have this factorization property.
For $R=[r,r]$
one can alternatively represent $\bar S$ as a triple sum with 
the Littlewood-Richardson weights
\be
\bar S_{\mu\nu}^{[r,r]} = \sum_{\mu'\nu'}
\sum_{\mu,\nu\subset \lambda\subset [r,r]} C^\lambda_{\mu\mu'}C^\lambda_{\nu\nu'}
B^{\lambda}_{\mu'\nu'}
\label{SCCB}
\ee
then $B^\lambda_{\mu'\nu'}$ will be factorized combinations of quantum numbers,
but the number of sums seem to grow further for $R=[r^s]$ with $s\geq 3$
\cite{M333}.

Specifics of symmetric representations is that for them the skew characters
$\chi^*_{[\lambda]/[\mu]} = \chi^*_{[\lambda-\mu]}$ {\it do} factorize,
$C^{[\lambda]}_{[\mu],[\mu']} = \delta_{\mu+\mu',\lambda}$
and also the sum over $\lambda$ is a one-fold sum -- what reduces the generic
triple sum (\ref{SCCB}) over Young diagrams to the ordinary $q$-hypergeometric
polynomial, though of a rather complicated Askey-Wilson type ${_4\phi_3}$
with the fixed hypergeometric argument $z=q^2$.


\section*{Acknowledgements}

This work was performed at the Institute for the Information Transmission Problems
with the support from the Russian Science Foundation, Grant No.14-50-00150.

\end{document}